Magnolia Manufacturing makes wing components for large aircraft. Kevin Choi is the production manager, responsible for manufacturing, and Michelle Michaels is the marketing manager. Both managers are paid a flat salary and are eligible for a bonus. The bonus is equal to 1 percent of their base salary for every 10 percent profit that exceeds a target. The maximum bonus is 5 percent of salary. Kevin’s base salary is $360,000 and Michelle’s is $420,000.

The target profit for this year is $8 million. Kevin has read about a new manufacturing technique that would increase annual profit by 20 percent. He is unsure whether to employ the new technique this year, wait, or not employ it at all. Using the new technique will not affect the target.

Required:

a. Suppose that profit without using the technique this year will be $8 million. By how much will Kevin’s and Michelle’s bonus change if Kevin decides to employ the new technique?

Bonus Change
Kevin's
Michelle's

b. Suppose that profit without using the technique this year will be $10.5 million. By how much will Kevin’s and Michelle’s bonus change if Kevin decides to employ the new technique?

Bonus Change
Kevin's
Michelle's
Answer #1

1. if current profit is 8 millions-

current profit target profit excess % excess % bonus bonus kevin bonus michelle
8 million 8 million 0 0 0 0 0

after new technique

estimated profit target profit excess % excess % bonus bonus kevin bonus michelle
9.6 million 8 million 1.6 million 20% 2% 360000*2% =7200 420000*2% = 8400

change in bonus-

Kevin = 7200-0 = 7200; Michelle = 8400-0 = 8400

2.

if current profit is 10.5 millions-

current profit target profit excess % excess % bonus bonus kevin bonus michelle
10.5 million 8 million 2.5 31.25% 3% 360000*3% = 10800 420000*3% = 12600

after new technique

estimated profit target profit excess % excess % bonus bonus kevin bonus michelle
12.6 million 8 million 4.6 million 57.5% 5% 360000*5% =18000 420000*5% = 21000

change in bonus-

Kevin = 18000-10800 = 7200; Michelle = 21000-12600 = 8400